If 5 apples in a barrel of 25 apples are rotten, what is the expected number of rotten apples in a random sample of 2 apples?
Answer by stanbon(48568)

If 5 apples in a barrel of 25 apples are rotten, what is the expected number of rotten apples in a random sample of 2 apples?
--------
The random variable values are 0,1,2
P(0 rotten)= [20C2]/[25C2]=0.63
P(1 rotten)= 5C1*20C1/25C2=0.33
P(2 rotten)= [5C2]/25C2=0.03
Expected number of rotten =0.63*0+0.33*1+0.03*2=0.39

Cheers,
Stan H.